Velma and Keota (V&K) is a partnership that owns a small company. It is considering two alternative investment opportunities. The first investment opportunity will have a four-year useful life, will cost $9,395.19, and will generate expected cash inflows of $2,900 per year. The second investment is expected to have a useful life of three years, will cost $7,473.58, and will generate expected cash inflows of $2,900 per year. Assume that V&K has the funds available to accept only one of the opportunities. (PV of $1 and PVA of $1) (Use appropriate factor(s) from the tables provided.)

Required

  1. Calculate the internal rate of return of each investment opportunity. (Do not round intermediate calculations.)

  2. Based on the internal rates of return, which opportunity should V&K select?
Answer #1

a. Internal rate of return = Cost / annual cash inflow

Find the value in cumulative present value factor

First alternative = 9395.19 / 2900 = 3.240 = 9%

Second alternative = 7473.58 / 2900 = 2.577 = 8%

b. First alternative should V&K select.